If the object has a uniform mass distribution a=6 m b=12 m c= 24 m find the center of gravity of the object
1
Expert's answer
2021-08-23T12:35:22-0400
"x=\frac{6}{2}=3\ m\\y=\frac{12}{2}=6\ m\\z=\frac{24}{2}=12\ m\\"